# Minimal Torsion Curves in Geometric Isogeny Classes

### v1

@@ 6.1. Preliminary Results
    $\operatorname{im}\rho_{E_{0},3^{\infty}}=9.36.0.7$ or $9.36.0.9$: As in case (2), the isogeny class 1734.k in the LMFDB shows we are free to assume $E_{0}$ has image 9.36.0.7. A Magma computation shows that there exists a cubic extension $L$ such that a twist $E_{0}^{t}$ of $E_{0}/L$ has an $L$-rational point of order 9 (say $Q$) and an independent 3-isogeny (say, with kernel generated by $R$). Then $\psi:E_{0}^{t}\rightarrow E_{1}=E_{0}^{t}/\langle R\rangle$ is a degree 3 isogeny, where $E_{1}$ has an $L$-rational cyclic 27-isogeny and $\psi(Q)\in E_{1}(F)$ is a point of order 9. Moreover, $\psi(Q)$ is in the kernel of the rational 27-isogeny. Thus the image of the 27-isogeny character $\chi$ associated to $E_{1}/L$ lands in $\{1,10,19\}$ and $E_{1}$ attains a point of order 27 in $\overline{L}^{\ker(\chi)}$, an extension of $L$ of degree dividing 3. Hence $E_{1}$ corresponds to a point on $X_{1}(27)$ of degree dividing 9. Since $d=2$, the divisibility conditions of Proposition 3.2 are best-possible for $k\geq 3$ with $j_{\text{min}}=j(E_{1})$. No elliptic curve in $\mathcal{E}$ with $j$-invariant in $\mathbb{Q}$ has a point of order 27 in this degree or lower.
@@
    $\operatorname{im}\rho_{E_{0},3^{\infty}}=9.36.0.8$: By Proposition 6.6, we have $3^{\max(0,2k-3)}\mid[F:\mathbb{Q}]$. A Magma computation shows that for $E_{0}/\mathbb{Q}$ with this image, there exists a cubic extension $L$ such that $E_{0}/L$ has an $L$-rational 9-isogeny and an independent 3-isogeny. As in case (3), there exists $E_{1}$ which is 3-isogenous to $E_{0}$ and gives a closed point on $X_{1}(27)$ of degree at most 27. Thus the divisibility conditions of Proposition 6.6 are best-possible for $k\geq 3$ with $j_{\text{min}}=j(E_{1})$. No elliptic curve in $\mathcal{E}$ with $j$-invariant in $\mathbb{Q}$ has a point of order 27 in this degree or lower.
@@ 7. Results for non-CM classes and $\ell=2$
Any non-CM $\mathbb{Q}$-curve defined over a number field of odd degree is geometrically isogenous to an elliptic curve with rational $j$-invariant, by work of Cremona and Najman [14, Theorem 2.7]. Thus the following can be viewed as a strengthening of [8, Proposition 4.1], which shows that if a non-CM $\mathbb{Q}$-curve has a point of order $2^{k}$ over a field of odd degree then $k\leq 4$. There exist non-CM elliptic curves over $\mathbb{Q}$ with a rational point of order $8$, so the following gives the best possible bound. The following proposition proves Theorem 1.1 in the case of $\ell=2$.
@@
The dual isogeny $\hat{\varphi}:E^{\prime}\rightarrow E$ is also cyclic of degree $2^{r}$, and so $E^{\prime}$ possesses an $F$-rational cyclic subgroup of order $2^{r}$. Rational subgroups are twist-invariant, so any elliptic curve $E_{0}/\mathbb{Q}$ with $j(E_{0})=j(E^{\prime})$ will possess an $F$-rational subgroup of order $2^{r}$, say generated by $Q$. It follows that $2^{r-1}Q$ is $F$-rational. Since $F$ has odd degree, it must be that $[\mathbb{Q}(2^{r-1}Q):\mathbb{Q}]=1$ or 3. If $r=1$, then it follows $\mathbb{Q}(\langle Q\rangle)=\mathbb{Q}(2^{r-1}Q)$. Otherwise, by [14, Proposition 3.6],
$[\mathbb{Q}(\langle Q\rangle):\mathbb{Q}(\langle 2Q\rangle)]\leq 2.$
Since $\mathbb{Q}(\langle Q\rangle)$ is contained if $F$, it must be of odd degree, and so $\mathbb{Q}(\langle Q\rangle)=\mathbb{Q}(\langle 2^{r-1}Q\rangle)=\mathbb{Q}(2^{r-1}Q)$. In either case, $\mathbb{Q}(\langle Q\rangle)$ is of degree 1 or 3. If $\mathbb{Q}(\langle Q\rangle)=\mathbb{Q}$, then $j(E)\in\mathbb{Q}$ and this contradicts [7, Theorem 3]. Thus $\mathbb{Q}(\langle Q\rangle)$ is of degree 3.
Next, we will show $r=1$. If not, then both $\langle 2^{r-2}Q\rangle$ and $2^{r-1}Q$ generate the same degree 3 extension of $\mathbb{Q}$. By Proposition 2.3, the elliptic curve $E_{0}$ has a closed point of degree 3 on $X_{1}(4)$ lying above a degree 3 point on $X_{1}(2)$. By the classification of 2-adic images due to Rouse and Zureick-Brown [33], this implies $E_{0}$ has 2-adic image X20b, X20a, or X20; see the data file associated to Corollaries 3.4 and 3.5 of [22]. These groups have RSZB labels 4.16.0.2, 8.16.0.3, and 4.8.0.2, respectively. Thus $\operatorname{ord}_{2}([\operatorname{GL}_{2}(\mathbb{Z}_{2}):\operatorname{im}\rho_{E_{0},2^{\infty}}])=3$ or 4. However, this implies the degree of $F$ is even by Proposition 3.2. Therefore $r=1$ and $Q$ has order $2$.
@@
Finally, we will prove the refined degree bounds for each $k\leq 3$. As above, suppose $x\in X_{1}(2^{k})$ is a point of odd degree associated to $E\in\mathcal{E}$ and $F\coloneqq\mathbb{Q}(x)$. We may assume there is an $F$-rational cyclic isogeny $\varphi:E\rightarrow E^{\prime}$ of degree $2^{r}$ with $j(E^{\prime})\in\mathbb{Q}$, where $E(F)$ has a point of order $2^{k}$. Let $E_{0}/\mathbb{Q}$ with $j(E_{0})=j(E^{\prime})$. By the second paragraph, if $E_{0}$ has a rational point of order 2, then $j(E)\in\mathbb{Q}$ and $E$ gives a closed point of degree 1 on $X_{1}(2)$ by [14, Proposition 3.2]. By [23, Proposition 4.6], the only odd degree closed points on $X_{1}(2^{k})$ associated to $E$ have degree 1. So it suffices to consider the case when $E_{0}$ has no rational point of order 2.
[…]
By replacing $E_{0}$ by a quadratic twist if necessary, we may assume the 2-adic image contains $-I$. Thus we may assume $\operatorname{im}\rho_{E_{0},2^{\infty}}=2.2.0.1$ or 4.8.0.2; see [33] for curves that minimally cover X2 and X20. In the first, there are no odd degree points on $X_{1}(8)$ associated to $\mathcal{E}$ by Proposition 3.2, so assume $\operatorname{im}\rho_{E_{0},2^{\infty}}=4.8.0.2$. By Corollary 3.6, results can be obtained by choosing a particular elliptic curve with this Galois image and computing the degrees of points on $X_{1}(8)$ for elliptic curves $2^{r}$-isogenous to $E_{0}$, where the $j$-invariants of the isogenous curves are roots of the modular polynomial $\Phi_{2^{r}}(X,j(E_{0}))$. A Magma computation shows that 2 divides the degree of a point on $X_{1}(8)$ associated to any elliptic curve $2^{r}$-isogenous to $E_{0}$ for $r\in\mathbb{Z}^{+}$, as desired. ∎
@@ 8. $\ell$-adic images of level $\ell$
If $E_{0}/\mathbb{Q}$ has no rational $\ell$-isogeny and $\operatorname{im}\rho_{E_{0},\ell}$ is not surjective, then $\operatorname{im}\rho_{E_{0},\ell^{\infty}}$ is the complete preimage of one of the following groups (see §2.2§2.2$\S$ 2.2§ 2.2). One may check that $\operatorname{ord}_{\ell}([\operatorname{GL}_{2}(\mathbb{Z}_{\ell}):\operatorname{im}\rho_{E_{0},\ell^{\infty}}])=1$. We will consider each case separately. We will see $j_{min}\not\in\mathbb{Q}$ in each case.
@@
  • 13S4, 5S4: The curve $E_{0}$ attains two independent $\ell$-isogenies in degree 6 with kernels $C_{1}$ and $C_{2}$. By Proposition 2.3, the curve $E_{0}/C_{1}$ attains a closed point on $X_{1}(\ell)$ in degree dividing $6\cdot\frac{\varphi(\ell)}{2}=3(\ell-1)$ and a closed point on $X_{1}(\ell^{2})$ in degree dividing $6\cdot\frac{\varphi(\ell^{2})}{2}=3\ell(\ell-1)$. The claim holds by Propositions 3.2 and 2.6 with $j_{min}=j(E_{0}/C_{1})$.
  • 7Ns, 7Ns.2.1, 7Ns.3.1, 5Ns, 5Ns.2.1, 3Ns: The curve $E_{0}$ picks up 2 independent $\ell$-isogenies in degree 2 with kernels $C_{1}$ and $C_{2}$. By Proposition 2.3, the curve $E_{0}/C_{1}$ attains a closed point on $X_{1}(\ell)$ in degree dividing $12\cdot\frac{\varphi(\ell)}{2}=\ell-$ 1 and a closed point on $X_{1}(\ell^{2})$ in degree dividing $2\cdot\frac{\varphi(\ell^{2})}{2}=\ell(\ell-1)$. The claim holds by Propositions 3.2 and 2.6 with $j_{min}=j(E_{0}/C_{1})$.
Now suppose $\ell=2$. If $\operatorname{im}\rho_{E_{0},2}=$ 2Cs, then $E_{0}$ has full 2-torsion over $\mathbb{Q}$. Hence it is isogenous over $\mathbb{Q}$ to an elliptic curve $E^{\prime}/\mathbb{Q}$ with a $\mathbb{Q}$-rational cyclic 4-isogeny. By Proposition 2.3, there is a degree 1 closed point on $X_{1}(4)$ associated to $E^{\prime}$, and the claim follows from Proposition 3.2 and the formula for $\deg(X_{1}(2^{k})\rightarrow X_{1}(4)$. If $\operatorname{im}\rho_{E_{0},2}=$ 2B or 2Cn, then $E_{0}$ has full 2-torsion over an extension $K$ of degree 2 or 3, respectively. There is $E^{\prime}/K$ with a $K$-rational cyclic 4-isogeny, and the claim follows as in the previous case. ∎
@@ 9. CM Elliptic Curves
Let $\mathcal{E}$ be a $\overline{\mathbb{Q}}$-isogeny class of CM elliptic curves. The endomorphism algebra $K=\text{End}(E)\otimes\mathbb{Q}$ is an isogeny invariant, so all elliptic curves in $\mathcal{E}$ have CM by an order in the imaginary quadratic field $K$. In fact, since any CM elliptic curve is isogenous to one with CM by the maximal order (see, for example, [9, Proposition 2.2]), the class $\mathcal{E}$ contains elliptic curves with CM by any possible order in $K$. In this section, we study the isogeny distance from a minimal torsion curve to an elliptic curve $E$ with CM by the full ring of integers in $K$, since $[\mathbb{Q}(j(E)):\mathbb{Q}]$ is minimal for $\mathcal{E}$. This builds on work of the first author and Clark [3, 4]. A key first step is to establish sharp lower bounds on the least degree of a point on $X_{1}(\ell^{k})$ associated to $\mathcal{E}$. These appear as Propositions 9.1, 9.2, and 9.3. Taken together they imply Theorem 1.4. Preliminary results about CM elliptic curves are summarized in Section 2.5.

### v2

@@ 6.1. Preliminary Results
    $\operatorname{im}\rho_{E_{0},3^{\infty}}=9.36.0.7$ or $9.36.0.9$: As in case (2), the isogeny class 1734.k in the LMFDB shows we are free to assume $E_{0}$ has image 9.36.0.7. A Magma computation shows that there exists a cubic extension $L$ such that a twist $E_{0}^{t}$ of $E_{0}/L$ has an $L$-rational point of order 9 (say $Q$) and an independent 3-isogeny (say, with kernel generated by $R$). Then $\psi:E_{0}^{t}\rightarrow E_{1}=E_{0}^{t}/\langle R\rangle$ is a degree 3 isogeny, where $E_{1}$ has an $L$-rational cyclic 27-isogeny and $\psi(Q)\in E_{1}(F)$ is a point of order 9. Moreover, $\psi(Q)$ is in the kernel of the rational 27-isogeny. Thus the image of the 27-isogeny character $\chi$ associated to $E_{1}/L$ lands in $\{1,10,19\}$ and $E_{1}$ attains a point of order 27 in $\overline{L}^{\ker(\chi)}$, an extension of $L$ of degree dividing 3. Hence $E_{1}$ corresponds to a point on $X_{1}(27)$ of degree dividing 9. Since $d=2$, the divisibility conditions of Proposition 3.2 are best-possible for $k\geq 3$ with $j_{min}=j(E_{1})$. No elliptic curve in $\mathcal{E}$ with $j$-invariant in $\mathbb{Q}$ has a point of order 27 in this degree or lower.
@@
    $\operatorname{im}\rho_{E_{0},3^{\infty}}=9.36.0.8$: By Proposition 6.6, we have $3^{\max(0,2k-3)}\mid[F:\mathbb{Q}]$. A Magma computation shows that for $E_{0}/\mathbb{Q}$ with this image, there exists a cubic extension $L$ such that $E_{0}/L$ has an $L$-rational 9-isogeny and an independent 3-isogeny. As in case (3), there exists $E_{1}$ which is 3-isogenous to $E_{0}$ and gives a closed point on $X_{1}(27)$ of degree at most 27. Thus the divisibility conditions of Proposition 6.6 are best-possible for $k\geq 3$ with $j_{min}=j(E_{1})$. No elliptic curve in $\mathcal{E}$ with $j$-invariant in $\mathbb{Q}$ has a point of order 27 in this degree or lower.
@@ 7. Results for non-CM classes and $\ell=2$
Any non-CM $\mathbb{Q}$-curve defined over a number field of odd degree is geometrically isogenous to an elliptic curve with rational $j$-invariant, by work of Cremona and Najman [CremonaNajmanQCurve, Theorem 2.7]. Thus the following can be viewed as a strengthening of [OddDegQCurve, Proposition 4.1], which shows that if a non-CM $\mathbb{Q}$-curve has a point of order $2^{k}$ over a field of odd degree then $k\leq 4$. There exist non-CM elliptic curves over $\mathbb{Q}$ with a rational point of order $8$, so the following gives the best possible bound. The following proposition proves Theorem 1.2 in the case of $\ell=2$.
@@
The dual isogeny $\hat{\varphi}:E^{\prime}\rightarrow E$ is also cyclic of degree $2^{r}$, and so $E^{\prime}$ possesses an $F$-rational cyclic subgroup of order $2^{r}$. Rational subgroups are twist-invariant, so any elliptic curve $E_{0}/\mathbb{Q}$ with $j(E_{0})=j(E^{\prime})$ will possess an $F$-rational subgroup of order $2^{r}$, say generated by $Q$. It follows that $2^{r-1}Q$ is $F$-rational. Since $F$ has odd degree, it must be that $\mathbb{Q}(2^{r-1}Q):\mathbb{Q}]=1$ or 3. If $r=1$, then it follows $\mathbb{Q}(\langle Q\rangle)=\mathbb{Q}(2^{r-1}Q)$. Otherwise, by [CremonaNajmanQCurve, Proposition 3.6],
$\mathbb{Q}(\langle Q\rangle):\mathbb{Q}(\langle 2Q\rangle)]\leq 2.$
Since $\mathbb{Q}(\langle Q\rangle)$ is contained if $F$, it must be of odd degree, and so $\mathbb{Q}(\langle Q\rangle)=\mathbb{Q}(\langle 2^{r-1}Q\rangle)=\mathbb{Q}(2^{r-1}Q)$. In either case, $\mathbb{Q}(\langle Q\rangle)$ is of degree 1 or 3. If $\mathbb{Q}(\langle Q\rangle)=\mathbb{Q}$, then $j(E)\in\mathbb{Q}$ and this contradicts [OddDeg, Theorem 3]. Thus $\mathbb{Q}(\langle Q\rangle)$ is of degree 3.
Next, we will show $r=1$. If not, then both $\langle 2^{r-2}Q\rangle$ and $2^{r-1}Q$ generate the same degree 3 extension of $\mathbb{Q}$. By Proposition 2.3, the elliptic curve $E_{0}$ has a closed point of degree 3 on $X_{1}(4)$ lying above a degree 3 point on $X_{1}(2)$. By the classification of 2-adic images due to Rouse and Zureick-Brown [RouseDZB], this implies $E_{0}$ has 2-adic image X20b, X20a, or X20; see the data file associated to Corollaries 3.4 and 3.5 of [GJLR]. These groups have RSZB labels 4.16.0.2, 8.16.0.3, and 4.8.0.2, respectively. Thus $\operatorname{ord}_{2}([\operatorname{GL}_{2}(\mathbb{Z}_{2}):\operatorname{im}\rho_{E_{0},2^{\infty}}])=3$ or 4. However, this implies the degree of $F$ is even by Proposition 3.2. Therefore $r=1$ and $Q$ has order $2$.
@@
Finally, we will prove the refined degree bounds for each $k\leq 3$. As above, suppose $x\in X_{1}(2^{k})$ is a point of odd degree associated to $E\in\mathcal{E}$ and $F\coloneqq\mathbb{Q}(x)$. We may assume there is an $F$-rational cyclic isogeny $\varphi:E\rightarrow E^{\prime}$ of degree $2^{r}$ with $j(E^{\prime})\in\mathbb{Q}$, where $E(F)$ has a point of order $2^{k}$. Let $E_{0}/\mathbb{Q}$ with $j(E_{0})=j(E^{\prime})$. By the second paragraph, if $E_{0}$ has a rational point of order 2, then $j(E)\in\mathbb{Q}$ and $E$ gives a closed point of degree 1 on $X_{1}(2)$ by [CremonaNajmanQCurve, Proposition 3.2]. By [GJNajman, Proposition 4.6], the only odd degree closed points on $X_{1}(2^{k})$ associated to $E$ have degree 1. So it suffices to consider the case when $E_{0}$ has no rational point of order 2.
The results for $X_{1}(2)$ are immediate, so suppose $2\leq k\leq 3$. By [OddDegQCurve, Lemma 6.3], the curve $E_{0}/\mathbb{Q}\in\mathcal{E}$ has full 2-torsion or a 4-isogeny defined over a number field of odd degree. Since $E_{0}(\mathbb{Q})$ has no point of order 2, these conditions occur if $E_{0}$ has $\operatorname{im}\rho_{E_{0},2}=\text{2Cn}$ or if $\operatorname{im}\rho_{E_{0},2^{\infty}}\in\{4.16.0.2,8.16.0.3,4.8.0.2\}$, respectively, as above. In either case, $E_{0}$ or an elliptic curve 2-isogenous to $E_{0}$ has a cyclic 4-isogeny defined over a cubic field. This gives a degree 3 closed point on $X_{1}(4)$ by Proposition 2.3. However, we will show no point on $X_{1}(8)$ associated to $\mathcal{E}$ has odd degree.
By replacing $E_{0}$ by a quadratic twist if necessary, we may assume the 2-adic image contains I-I. Thus we may assume $\operatorname{im}\rho_{E_{0},2^{\infty}}=2.2.0.1$ or 4.8.0.2; see [RouseDZB] for curves that minimally cover X2 and X20. In the first, there are no odd degree points on $X_{1}(8)$ associated to $\mathcal{E}$ by Proposition 3.2, so assume $\operatorname{im}\rho_{E_{0},2^{\infty}}=4.8.0.2$. By Corollary 3.6, results can be obtained by choosing a particular elliptic curve with this Galois image and computing the degrees of points on $X_{1}(8)$ for elliptic curves $2^{r}$-isogenous to $E_{0}$, where the $j$-invariants of the isogenous curves are roots of the modular polynomial $\Phi_{2^{r}}(X,j(E_{0}))$. A Magma computation shows that 2 divides the degree of a point on $X_{1}(8)$ associated to any elliptic curve $2^{r}$-isogenous to $E_{0}$ for $r\in\mathbb{Z}^{+}$, as desired. ∎
@@ 8. $\ell$-adic images of level $\ell$
If $E_{0}/\mathbb{Q}$ has no rational $\ell$-isogeny and $\operatorname{im}\rho_{E_{0},\ell}$ is not surjective, then $\operatorname{im}\rho_{E_{0},\ell^{\infty}}$ is the complete preimage of one of the following groups (see $\S 2.2$). One may check that $\operatorname{ord}_{\ell}([\operatorname{GL}_{2}(\mathbb{Z}_{\ell}):\operatorname{im}\rho_{E_{0},\ell^{\infty}}])=1$. We will consider each case separately. We will see $j_{min}\not\in\mathbb{Q}$ in each case.
@@
  • 13S4, 5S4: A Magma computation shows the curve $E_{0}$ attains two independent $\ell$-isogenies in degree 6 with kernels $C_{1}$ and $C_{2}$. By Proposition 2.3, the curve $E_{0}/C_{1}$ attains a closed point on $X_{1}(\ell)$ in degree dividing $6\cdot\frac{\varphi(\ell)}{2}=3(\ell-1)$ and a closed point on $X_{1}(\ell^{2})$ in degree dividing $6\cdot\frac{\varphi(\ell^{2})}{2}=3\ell(\ell-1)$. The claim holds by Propositions 3.2 and 2.6 with $j_{min}=j(E_{0}/C_{1})$.
  • 7Ns, 7Ns.2.1, 7Ns.3.1, 5Ns, 5Ns.2.1, 3Ns: The curve $E_{0}$ picks up 2 independent $\ell$-isogenies in degree 2 with kernels $C_{1}$ and $C_{2}$. By Proposition 2.3, the curve $E_{0}/C_{1}$ attains a closed point on $X_{1}(\ell)$ in degree dividing $2\cdot\frac{\varphi(\ell)}{2}=\ell-1$ and a closed point on $X_{1}(\ell^{2})$ in degree dividing $2\cdot\frac{\varphi(\ell^{2})}{2}=\ell(\ell-1)$. The claim holds by Propositions 3.2 and 2.6 with $j_{min}=j(E_{0}/C_{1})$.
Now suppose $\ell=2$. If $\operatorname{im}\rho_{E_{0},2}=$ 2Cs, then $E_{0}$ has full 2-torsion over $\mathbb{Q}$. Hence it is isogenous over $\mathbb{Q}$ to an elliptic curve $E^{\prime}/\mathbb{Q}$ with a $\mathbb{Q}$-rational cyclic 4-isogeny. By Proposition 2.3, there is a degree 1 closed point on $X_{1}(4)$ associated to $E^{\prime}$, and the claim follows from Proposition 3.2 and the formula for $\deg(X_{1}(2^{k})\rightarrow X_{1}(4))$. If $\operatorname{im}\rho_{E_{0},2}=$ 2B or 2Cn, then $E_{0}$ has full 2-torsion over an extension $K$ of degree 2 or 3, respectively. There is $E^{\prime}/K$ with a $K$-rational cyclic 4-isogeny, and the claim follows as in the previous case. ∎
@@
The next result shows that working with curves in a non-CM rational geometric isogeny class $\mathcal{E}$ can yield points of strictly lower degree than those associated with any rational non-CM $j$-invariant. This relies on recent work of Furio [Furio2024], as applied in work of the first author with Ejder [BourdonEjder].
Corollary 8.3.
The least degree of a point on $X_{1}(49)$ associated to any non-CM rational $\overline{\mathbb{Q}}$-isogeny class $\mathcal{E}$ is at most 42, whereas the least degree of a non-CM point $x\in X_{1}(49)$ with $j(x)\in\mathbb{Q}$ is 49.
Proof.
The first claim follows from the previous proposition: Consider, for example, an elliptic curve over $\mathbb{Q}$ with mod 7 image 7Ns. For the second claim, note that all known non-surjective 7-adic images for non-CM elliptic curves $E/\mathbb{Q}$ have level 7; see [RouseDZB]. Thus any point on $X_{1}(49)$ associated to such an $E$ has degree at least 49. Equality holds for any elliptic curve over $\mathbb{Q}$ with a rational point of order 7. If $E/\mathbb{Q}$ has surjective 7-adic image, then $E$ gives a single closed point on $X_{1}(49)$ of degree 1176. By [RSZ21, Theorem 1.1.6] and [FurioLombardo25, Theorem 1.4], other groups which could occur as 7-adic images for non-CM elliptic curves over $\mathbb{Q}$ would give rise to rational points on the modular curves with RSZB label 49.147.9.1 or 49.196.9.1. In the first case, such an $E/\mathbb{Q}$ would have mod 7 image landing in the normalizer of a non-split Cartan subgroup. Possible 7-adic images for $E$ are given in [Furio2024, Theorem 1.9], and the last paragraph of the proof of Theorem 6 in [BourdonEjder] shows $E$ gives a point on $X_{1}(49)$ of degree at least 168. Finally, if $E/\mathbb{Q}$ corresponds to a rational point on the modular curve labeled 49.196.9.1, then $\operatorname{im}\rho_{E,7^{\infty}}=49.196.9.1$ by [BourdonEjder, Corollary 3]. Such an image gives points on $X_{1}(49)$ of degree at least 294. ∎
@@ 9. CM Elliptic Curves
Let $\mathcal{E}$ be a $\overline{\mathbb{Q}}$-isogeny class of CM elliptic curves. The endomorphism algebra $K=\text{End}(E)\otimes\mathbb{Q}$ is an isogeny invariant, so all elliptic curves in $\mathcal{E}$ have CM by an order in the imaginary quadratic field $K$. In fact, since any CM elliptic curve is isogenous to one with CM by the maximal order (see, for example, [BP, Proposition 2.2]), the class $\mathcal{E}$ contains elliptic curves with CM by any possible order in $K$. In this section, we study the isogeny distance from a minimal torsion curve to an elliptic curve $E$ with CM by the full ring of integers in $K$, since $\mathbb{Q}(j(E)):\mathbb{Q}]$ is minimal for $\mathcal{E}$. This builds on work of the first author and Clark [BC1, BC2]. A key first step is to establish sharp lower bounds on the least degree of a point on $X_{1}(\ell^{k})$ associated to $\mathcal{E}$. These appear as Propositions 9.1, 9.2, and 9.3. Taken together they imply Theorem 1.1. Preliminary results about CM elliptic curves are summarized in Section 2.5.